OpenStudy (anonymous):

How many grams of water can be cooled from 35ºC to 20ºC by the evaporation of 60g water? (The heat of evaporation is 2.4 kJ/g in this temperature range and the specific heat of water is 4.18 J/ g-K.)

OpenStudy (rina.r):

It takes energy to evaporate water, if this energy is taken from a pool of water, and it's cooled from 35 to 23C, then how big a pool was it. Calculate how much energy it takes to evaporate 39 g: H=H_vap*m=2.4*39=93.6 KJ Now use: H=m*H_s*delta_T=93.6KJ and solve for m m=(93.6*10^3)/(12*4.18)=1860 g Notes: change in temp should be in Kelvin, but a degree in C = a degree in K watch units...H_vap is in KJ but H_s is in J... that's why there is a 10^3 in calculation of m http://answers.yahoo.com/question/index?qid=20100119145457AA5UwN2